OpenStudy (anonymous):

Write a polynomial function in standard form wtih zeros at 5, -4, and -3? a. f(x)= x^3 - 60x^2 + 2x - 23 b. f(x)= x^3 + 2x^2 - 23x + 7 c. f(x)= x^3 - 17x^2 - 420x +7 d. f(x)= x^3 + 2x^2 + 23x - 60

sam (.sam.):

roots are 5, -4, and -3 x=5 , x=-4 , x=-3 x-5=0 , x+4=0 , x+3=0 (x-5)(x+4)(x+3)=0 expand and simplify

OpenStudy (anonymous):

Thanks :) I copied the wrong answer list but now i know how to do it thanks :D